OpenStudy (anonymous):

What is the solution of the system of equations? y = –3x + 8 y = –5x – 2 A.(23, –5) B.(–5, 23) C.(–1.25, 11.75) D.(3, –17)

OpenStudy (photon336):

first we need to eliminate one of the variables. we can multiply the top by -5 and the bottom by +3 like this -5(y = -3x+8) = -5y = 15x-40 3(y = -5x-2) = 3y = -15x-6 we do this to eliminate one of the variables let's say y. and we add both of them together: we notice that x cancels out when we add the two. -5y = 15x-40 3y = -15x-6 -2y = -46 -2 -2 y = 23 plug this back into the original equation: 23 = -3x+8 -8 -8 15 = -3x -3 -3 x = -5 y = 23 let's plug these both into the second equation and see what we get: y = -5x-2 y = -5(-5)-2 -5(-5) = 25 -2 = 23 23 = 23 that's how we know we got it correct so ( y = 23, and x = -5)